OpenStudy (anonymous):

As the mass and distance between two objects decrease, so does the gravitational force exerted between them. Answer True False

OpenStudy (ivancsc1996):

If both the mass and distace decrease at the same rate, True

OpenStudy (ivancsc1996):

False sorry

OpenStudy (anonymous):

\[F = \frac{ GMm }{ d ^{2} }\] If the mass of either object decreases the force decreases. If the distance decreases the force increases. If both masses as well as the distance decrease at the same rate: Assume the masses and distance half \[F= \frac{ G \frac{ M }{ 2 } \frac{ m }{ 2 }}{ \left( \frac{ d }{ 2 } \right)^{2} }\] \[F = \frac{ \frac{ GMm }{ 4 }}{ \left( \frac{ d ^{2} }{ 4 } \right) }\] \[F = \frac{ GMm }{ 4 } \div \frac{ d ^{2} }{ 4 }\] \[F = \frac{ GMm }{ 4 } \times \frac{ 4 }{ d ^{2} }\] \[F = \frac{ GMm }{ d ^{2} }\] Therefore if both the masses and the distance decrease at the same rate there will be no change in force

OpenStudy (priths):

How can the mass between two objects decrease ? I think its false.